# ComparisonQA: Evaluating Factuality Robustness of LLMs Through Knowledge Frequency Control and Uncertainty

### v1

@@ Abstract
The rapid development of LLMs has sparked extensive research into their factual knowledge. Current works claim that LLMs fall short on questions requiring less frequent knowledge. However, their proof is incomplete since they only study the influence of entity frequency, which can not fully represent knowledge frequency. So we introduce ComparisonQA benchmark, containing 283K abstract questions, each instantiated by a pair of high-frequency and low-frequency entities. It ensures a controllable comparison because the difference of knowledge frequency between such a pair is only related to entity frequency. In addition, to avoid possible semantic shortcuts, which is a severe problem of current LLMs study, we design a two-round method for knowledge robustness measurement utilizing both correctness and uncertainty. Experiments reveal that LLMs exhibit particularly low robustness regarding low-frequency knowledge, and GPT-4o is even the worst under this measurement. Besides, we introduce an automatic method to filter out questions with low-quality and shortcuts to form ComparisonQA-Hard. We find that uncertainty effectively identifies such questions while maintaining the data size.  [^1] https://github.com/HKUST-KnowComp/ComparisonQA
@@
Qing Zong, Zhaowei Wang, Tianshi Zheng, Xiyu Ren, Yangqiu Song Department of Computer Science and Engineering, HKUST {qzong, zwanggy, tzhengad, xrenaf, yqsong}@cse.ust.hk
@@ 1 Introduction
The rapid advancement of large language models (LLMs) has promoted a lot of study on their factual knowledge and reasoning ability, and many factuality QA benchmarks Wei et al. (2024a, b); Hendrycks et al. (2021); Kwiatkowski et al. (2019) have been introduced accordingly.
@@
Sun et al. (2024) and Mallen et al. (2023) find that LLM’s performance drops as the frequency of the knowledge required to solve the question decreases. They compare LLM’s performance on questions about entities with different frequencies. However, we argue that such comparisons are incomplete and unfair since entity frequency can not fully represent the knowledge frequency. For example, consider the two questions: When was [Einstein] born and What was [Einstein]’s favorite fruit. Although both questions involve well-known entities, they have quite different difficulties, since Einstein’s birthday can be found in many places while his favorite fruit is rarely mentioned.
@@
To tackle the issue, we introduce our ComparisonQA benchmark. Pairs of high-frequency and low-frequency entities share the same question, as shown in Figure 1. The shared abstract question, with a hypernym to represent the two specific entities, guarantees that the difference of the required knowledge between such a pair is only related to the entities. This allows for a controllable comparison between high-frequency and low-frequency entities. ComparisonQA is a large scale dataset containing 283K such question pairs. It is generated through an automatic pipeline from raw knowledge, ensuring both diversity and scalability. Through this, we completely prove that LLM performance is closely related to the frequency of the knowledge required.
@@
In our benchmark, we use the multiple-choice format Hendrycks et al. (2021); Geva et al. (2021) for a more robust and accurate evaluation. But semantic shortcuts Geirhos et al. (2020) between questions and options may help LLMs to guess the answer, which is a common but severe problem recently. Thus, we design a two-round method to evaluate LLMs’ robust knowledge through correctness and uncertainty. During experiments, we found that LLMs have very low robustness, especially on questions requiring low-frequency knowledge, where the powerful GPT-4o even performed the worst among all LLMs we tested.
@@
We also provide a subset called ComparisonQA-Hard for future study. It contains 81K difficult low-frequency questions with high quality and no semantic shortcuts. We propose an automatic method that also uses correctness and uncertainty to filter out questions with low-quality and shortcuts. Experiments show that uncertainty is more effective in filtering these questions while maintaining the size of the benchmark, compared to the recent method used in benchmarks like SimpleQA Wei et al. (2024a), which only collects questions adversarially against LLMs’ responses.
@@
In summary, we have three main contributions: (1) [Resource] We introduce ComparisonQA benchmark, where a pair of entities share the same abstract question. It enables a more reasonable proof that LLMs perform worse when the required knowledge is less frequent. (§3)
@@
(2) [Method] We propose a two-round method using correctness and uncertainty for LLMs’ robust knowledge measurement. [Finding] LLMs can not stand such a test, especially on the low-frequency knowledge, where GPT-4o even performs the worst among all tested LLMs. (§4)
@@
(3) [Resource] We introduce ComparisonQA-Hard benchmark containing only hard and low-frequency questions of high quality and no shortcuts. [Finding] When doing the automatic filtering, uncertainty is more helpful to find questions with low quality and shortcuts. (§5)
@@ 2.1 Benchmarking LLMs’ Factuality
The factuality evaluation of LLMs has recently attracted significant attention Wang et al. (2024c). Some factuality benchmarks require open-ended generation by LLMs, such as SimpleQA Wei et al. (2024a), FreshQA Vu et al. (2024), SelfAware Yin et al. (2023), CLR-Fact Zheng et al. (2024a), and HaluEval Li et al. (2023). Such evaluations either rely heavily on expert annotation, or utilize automatic answer matching that sacrifices evaluation accuracy Min et al. (2023); Chern et al. (2023); Wang et al. (2024b). Other benchmarks adopt the format of Yes-or-No questions Geva et al. (2021); Zhang et al. (2024) or multiple-choice questions (MCQ) Hendrycks et al. (2021); Wang et al. (2024a). These formats allow model responses to be easily parsed and compared with gold labels, enabling solid yet efficient evaluations.
Figure 2: An overview of our benchmark curation pipeline. It contains three parts. Through the first two parts, (1) Entity Pairs Extraction from DBpedia and (2) Abstract Question Generation, we can get the whole ComparisonQA. And through the third part, (3) Simple Low-Quality Question Filtering, we can get a harder subset, containing only difficult low-frequency questions with high quality and no semantic shortcut.
@@ 2.2 Long-Tail Knowledge
Long-tail knowledge Wei et al. (2024b); Chen et al. (2023) is an important aspect of factuality. Kumar et al. (2024) proposes an automatic approach to generate questions for tail entities. Kandpal et al. (2023) find LLMs struggle to learn long-tail knowledge. Other works study the influence of knowledge frequency: Mallen et al. (2023) introduced PopQA, a long-tail benchmark, and found that models’ performance will change with the frequency of entities in the questions. Sun et al. (2024) also proves this by constructing questions around head, torso, and tail entities. However, questions in these benchmarks are all in the form of open-ended generation, which can not be easily evaluated. They also depend on limited number of templates to produce QA questions from knowledge graphs, which will significantly harm the diversity of the benchmarks. More importantly, they only consider the frequency of the entities involved but ignore the entire required knowledge.
@@ 2.3 Abstraction Knowledge
Existing works have studied various aspects of abstraction, for example, entity abstraction Wu et al. (2012); Song et al. (2015), event abstraction Wang et al. (2024e, d), and conceptual abstraction Han et al. (2024). Abstraction has been proven beneficial to downstream tasks like commonsense reasoning, numerical reasoning, and logical reasoning Zhou et al. (2024); Hong et al. (2024). In this paper, we control question difficulty by sharing the same abstraction form between a pair of entities (§3.1).
@@ 3.1 Question Formulation
For accurate evaluation, our questions are in the form of multiple-choice. Although recently there are several generative QA benchmarks, like SimpleQA Wei et al. (2024a), and also some automatic methods Zheng et al. (2023) to evaluate generated answers. There are still significant limitations in such generative QAs since the answer should be single and indisputable. Questions like What is the primary focus or intention behind Civil Procedure Rules cannot be included since there are many ways to answer this question. However, multiple-choice questions do not have such limitations.
Each piece of data in ComparisonQA, shown in Figure 1, contains an abstract question shared by two entities having the same hypernym. One entity, having many relationships in DBpedia Auer et al. (2007), which will be introduced next, is the high-frequency entity, and the other, with only a few relationships, is the low-frequency one. The question will have different correct answers for the two entities, respectively. Such kind of data form can help us make more detailed and controllable comparisons, and that is why we call it ComparisonQA. The whole benchmark, containing 283K such question pairs, is constructed through a fully automated pipeline, which is cheap and scalable.
@@ 3.2 Curation Pipeline
In this section, we discuss the data curation pipeline for our dataset. As shown in Figure 2, the pipeline contains three parts: (1) Entity Pairs Extraction from DBpedia, (2) Abstract Question Generation, and (3) Simple Low-Quality Question Filtering. Here we will introduce the first two parts, used to build ComparisonQA, and leave the third part, used to build ComparisonQA-Hard, for §5.
@@ 3.2.1 Entity Pairs Extraction from DBpedia
Following Sun et al. (2024), an entity’s frequency is defined by its number of relationships in DBpedia. High-frequency and low-frequency entities, whose cumulative relationships account for the first and last $1/3$ of all sorted entities, are those with relationships above 185 and below 107 separately, with experiments shown in Appendix B.
In practice, we first get all the hypernyms in DBpedia and classify the entities belonging to them into high-frequent and low-frequent based on these two relationship counts. Then, we map the entities one-by-one to get entity pairs. The reason behind this is that not every pair of entities can have a shared abstract question easily. For example, it’s hard to write a shared question for Einstein and Apple even though they are all high frequency entities. But it’s easy for two specific universities to share a same abstract question. In order for our entity pairs to produce high quality abstract questions, we ensure that the two entities have the same hypernym. This allows them to have similar descriptions, making it easier to generate an abstract question.
@@ 3.2.2 Abstract Question Generation
After acquiring the entity pairs, we adopt a multi-step curation pipeline to generate high-quality abstract multiple-choice questions: (1) Generate an abstract question without options according to the DBpedia descriptions of paired entities. (2) Separately generate the corresponding answers based on the descriptions, with length control to alleviate bias among candidate answers. (3) Generate distractors for both entities with length control. Compared to randomly selected distractors, LLM-selected distractors generation have demonstrated its effectiveness for the high relevance between distractors and designated choices Zheng et al. (2024b). (4) Formulate the final multiple-choice question using the four answer candidates above. (5) Proofread the question according to the standards below.
@@ 3.3 Expert Verification
We enlist the help of three postgraduate students, each with extensive experience in NLP research, to validate the quality of these generated questions through a sample of 200 question pairs. The quality of each pair is decided by majority voting, with standard shown above. Results show that their total agreement (all 3 experts agree with the quality) is 88.0%. And 95.5% of the abstract questions are considered correct and of high quality both for the high frequency entity and the low frequency entity, demonstrating the reliability of our benchmark.
@@ 3.4 Main Evaluations
We experiment with a selection of LLMs on our curated ComparisonQA benchmark to investigate their performance on high-frequency questions and low-frequency questions and also the difference between them.
@@ 3.4.1 Experiment Setup
Metric: We calculate the Uncertainty, Accuracy, and Macro F1-score between predictions and ground truth labels to evaluate the models. The calculation of uncertainty is explained in Appendix D.
@@ 3.4.2 Results and Analysis
Evaluation results are reported in Table 1. Our observations include: (1) Huge drop in performance from high-freq to low-freq: All models suffer a performance decrease from high-frequency questions to low-frequency questions in all three settings. For instance, the accuracy and Macro F1-score of Llama-3-8B drop up to about 14 points in the few-shot setting. Proprietary models like GPT-4o are no exception. These all prove that LLM’s performance is closely related to the frequency of the knowledge in the corpus. (2) Increased Uncertainty from high-freq to low-freq: Similarly, the uncertainty of LLMs all increase from high-frequency questions to low-frequency questions. For example, the uncertainty difference of Gemma-2-9B is up to about 86 points when using zero-shot. The uncertainty in the zero-shot setting is generally higher than in the few-shot setting, and the difference is also more pronounced. We think it may be because LLMs find some familiar examples in the few-shot setting, which decreases their uncertainty. In spite of this, the difference is still clear between high-freq and low-freq. These all prove that LLMs not only perform better but also are more confident about high-frequency knowledge. (3) Few-shot helps a lot only for LLMs without instruction-tuning: Most non-instruction-tuned LLMs show a huge improvement in performance from zero-shot to few-shot, but performance is similar for those instruction-tuned LLMs. This could be because the few-shot examples only teach LLMs how to do multiple-choice questions, while those instruction-tuned ones have already learned. (4) CoT lowers uncertainty but does not aid performance: It’s obvious that after the CoT inference, LLMs are more sure about their answers. However, results show that their performance does not have a big difference, which means CoT can not help with such factuality questions.
@@
Llama-3 8B 75.57 61.00 68.29 14.57 66.09 (-9.48)0 39.15 (-21.86) 52.62 (-15.67) ↓ 26.95 (+12.38)
Llama-3-Instruct 8B 79.94 67.98 73.96 ↓ 11.96 69.04 (-10.91) 43.50 (-24.49) 56.27 (-17.69) ↓ 25.54 (+13.58)
Llama-3.1 8B 74.91 62.00 68.46 ↓ 12.91 66.27 (-8.64)0 40.35 (-21.65) 53.31 (-15.15) ↓ 25.92 (+13.01)
Llama-3.1-Instruct 8B 79.74 69.09 74.42 ↓ 10.65 72.89 (-6.85)0 43.91 (-25.19) 58.40 (-16.02) 28.98 (+18.34)
Llama-3.2 3B 68.89 57.26 63.08 ↓ 11.63 59.15 (-9.73)0 38.19 (-19.07) 48.67 (-14.41) ↓ 20.96 (+9.33)0
Llama-3.2-Instruct 3B 71.43 62.12 66.78 09.32 64.14 (-7.30)0 40.19 (-21.93) 52.17 (-14.62) ↓ 22.68 (+14.63)
@@
Gemma-2 9B 80.10 68.36 74.23 ↓ 11.74 71.26 (-8.84)0 46.98 (-21.38) 59.12 (-15.11) ↓ 24.28 (+12.54)
Phi-3.5-mini 4B 73.68 67.46 70.57 06.22 67.08 (-6.60)0 44.39 (-23.06) 55.74 (-14.84) ↓ 22.69 (+16.47)
Falcon2 11B 77.11 65.92 71.52 ↓ 11.19 64.27 (-12.84) 43.37 (-22.55) 53.82 (-17.70) ↓ 20.90 (+9.71)0
Mistral-v0.3 7B 75.55 62.88 69.22 ↓ 12.67 68.75 (-6.80)0 39.81 (-23.07) 54.28 (-14.94) ↓ 28.94 (+16.27)
Mistral-v0.3-Instruct 7B 74.46 65.51 69.99 08.95 68.25 (-6.21)0 41.19 (-24.32) 54.72 (-15.27) ↓ 27.06 (+18.10)

### v2

@@ Abstract
The rapid development of LLMs has sparked extensive research into their factual knowledge. Current works find that LLMs fall short on questions around low-frequency entities. However, such proofs are unreliable since the questions can differ not only in entity frequency but also in difficulty themselves. So we introduce ComparisonQA benchmark, containing 283K abstract questions, each instantiated by a pair of high-frequency and low-frequency entities. It ensures a controllable comparison to study the role of knowledge frequency in the performance of LLMs. Because the difference between such a pair is only the entity with different frequencies. In addition, we use both correctness and uncertainty to develop a two-round method to evaluate LLMs’ knowledge robustness. It aims to avoid possible semantic shortcuts which is a serious problem of current QA study. Experiments reveal that LLMs, including GPT-4o, exhibit particularly low robustness regarding low-frequency knowledge. Besides, we find that uncertainty can be used to effectively identify high-quality and shortcut-free questions while maintaining the data size. Based on this, we propose an automatic method to select such questions to form a subset called ComparisonQA-Hard, containing only hard low-frequency questions.  [^1] https://github.com/HKUST-KnowComp/ComparisonQA
@@
Qing Zong,   Zhaowei Wang,   Tianshi Zheng, Xiyu Ren,   Yangqiu Song Department of Computer Science and Engineering, HKUST {qzong, yqsong}@cse.ust.hk
@@ 1 Introduction
The rapid advancement of large language models (LLMs) has promoted a lot of study on their factual knowledge and reasoning ability Wei et al. (2024a, b); Hendrycks et al. (2021).
@@
Sun et al. (2024) and Mallen et al. (2023) compare LLMs’ performance on questions around entities with different frequencies. LLMs are found struggle to handle tail knowledge. However, the questions they study are different and can vary in difficulty levels, not only in entity frequencies. As Allen-Zhu and Li (2023) emphasized, we need a more “controlled, synthetic experiment that confirms the weakness of LLMs” nowadays. Therefore, existing comparisons are not enough since they can not guarantee that the low frequency is the only cause of LLMs’ poor performance.
@@
To tackle the issue, we introduce our ComparisonQA benchmark. Pairs of high-frequency and low-frequency entities share the same question, as shown in Figure 1. The shared abstract question, with a hypernym to represent the two specific entities, guarantees that the difference between such a pair is only the entity. This allows for a controllable comparison between high-frequency and low-frequency entities. ComparisonQA is a large scale dataset containing 283K such question pairs. It is generated through an automatic pipeline base on raw knowledge base, ensuring both diversity and scalability. Through this benchmark, we can completely compare LLM’s performance on different knowledge frequencies.
@@
For a more robust and accurate evaluation, we use the multiple-choice format Hendrycks et al. (2021); Geva et al. (2021) in our benchmark. But semantic shortcuts Geirhos et al. (2020) between questions and options may help LLMs to guess the answer, which is also a common but severe problem recently. Thus, we further design a two-round method using both correctness and uncertainty to evaluate LLMs’ knowledge. During experiments, we found that LLMs have very poor robustness, especially on low-frequency knowledge, where even the powerful GPT-4o also performs badly.
@@
Recent benchmarks, like SimpleQA Wei et al. (2024a), ensure their difficulty by collecting questions adversarially against LLMs’ responses. But relying only on accuracy, they ignore the quality of their questions, and the difficulty is closely related to the models they use. Fortunately, our experiments find that uncertainty is also an effective tool in selecting both high-quality and shortcut-free questions while maintaining the benchmark size. Combining accuracy and uncertainty, we propose a new flexible method to select our subset called ComparisonQA-Hard for future study. It contains 81K difficult low-frequency questions with high-quality and no semantic shortcuts.
In summary, we have three main contributions:
@@
(1) [Resource] We introduce ComparisonQA benchmark, where a pair of entities share the same abstract question. It enables a more controllable and reasonable proof that LLMs perform worse when the required knowledge is less frequent. (§3)
@@
(2) [Method] We design a two-round method using correctness and uncertainty to evaluate LLMs’ robust knowledge. [Finding] LLMs can not stand such a test, especially on low-frequency knowledge, where even GPT-4o performs badly. (§4)
@@
(3) [Finding] Uncertainty is more helpful to find questions with high quality. [Resource] Through this, we select ComparisonQA-Hard benchmark containing only hard and low-frequency questions of high quality and no shortcuts. (§5)
@@ 2.1 Benchmarking LLMs’ Factuality
The factuality evaluation of LLMs has recently attracted significant attention. Some factuality benchmarks require open-ended generation by LLMs, such as SimpleQA Wei et al. (2024a), SelfAware Yin et al. (2023), CLR-Fact Zheng et al. (2024b), and HaluEval Li et al. (2023). Such evaluations either rely heavily on expert annotation, or utilize automatic answer matching that sacrifices evaluation accuracy Min et al. (2023); Chern et al. (2023); Wang et al. (2024b). Other benchmarks adopt the format of Yes-or-No questions Geva et al. (2021); Zhang et al. (2024) or multiple-choice questions (MCQ) Hendrycks et al. (2021); Wang et al. (2024a); Zheng et al. (2024a). These formats allow model responses to be easily parsed and compared with gold labels, enabling solid yet efficient evaluations.
Figure 2: An overview of our benchmark curation pipeline. It contains three parts. Through the first two parts, (1) Entity Pairs Extraction from DBpedia and (2) Abstract Question Generation, we can get the whole ComparisonQA. And through the third part, (3) Hard High-Quality Question Selecting, we can get a harder subset, containing only difficult low-frequency questions with high quality and no semantic shortcut.
@@ 2.2 Long-Tail Knowledge
Long-tail knowledge Wei et al. (2024b); Chen et al. (2023) is an important aspect of factuality. Kumar et al. (2024) proposes an automatic approach to generate questions for tail entities. Kandpal et al. (2023) find LLMs struggle to learn long-tail knowledge. Other works study the influence of knowledge frequency: Mallen et al. (2023) introduced PopQA, a long-tail benchmark, and found that models’ performance will change with the frequency of entities in the questions. Sun et al. (2024) also proves this by constructing questions around head, torso, and tail entities. However, questions in these benchmarks are all in the form of open-ended generation, which can not be easily evaluated. They also depend on limited number of templates to produce QA questions from knowledge graphs, which will significantly harm the diversity of the benchmarks. Most importantly, the questions are different, so they may vary in difficulty levels, and thus can not provide fair comparisons.
@@ 2.3 Abstraction Knowledge
Existing works have studied various aspects of abstraction, for example, entity abstraction Wu et al. (2012); Song et al. (2015); Xu et al. (2023), event abstraction Wang et al. (2024d, c), and conceptual abstraction Han et al. (2024). Abstraction has been shown to be beneficial for downstream tasks like commonsense reasoning, numerical reasoning, and logical reasoning Zhou et al. (2024); Hong et al. (2024). In this paper, we control question difficulty by sharing the same abstraction form between a pair of entities.
@@ 3.1 Question Formulation
For an accurate evaluation, our questions are in the form of multiple choice. Although recently there are several generative QA benchmarks, like SimpleQA Wei et al. (2024a), and also some automatic methods Zheng et al. (2023) to evaluate generated answers. There are still significant limitations in such generative QAs since the answer should be single and indisputable. Questions like What is the primary focus or intention behind Civil Procedure Rules cannot be included since there are many ways to answer this question. However, multiple-choice questions do not have such limitations.
Each piece of data in ComparisonQA, shown in Figure 1, contains an abstract question shared by two entities having the same hypernym. One entity, having many relationships in DBpedia Auer et al. (2007), which will be introduced next, is the high-frequency entity, and the other, with only a few relationships, is the low-frequency one. The question will have different answers for the two entities, respectively. Such data form can ensure detailed and controllable comparisons between entities with different frequencies, and that is why we call it ComparisonQA. Containing 283K such question pairs, the benchmark is constructed through a fully automated pipeline, which is cheap and scalable.
@@ 3.2 Curation Pipeline
In this section, we discuss the data curation pipeline for our dataset. As shown in Figure 2, the pipeline contains three parts: (1) Entity Pairs Extraction from DBpedia, (2) Abstract Question Generation, and (3) Hard High-Quality Question Filtering. Here we will introduce the first two parts, used to build ComparisonQA, and leave the third part, used to build ComparisonQA-Hard, for §5.
@@ 3.2.1 Entity Pairs Extraction from DBpedia
Following Sun et al. (2024), an entity’s frequency is defined by its number of relationships in DBpedia. High-frequency entities are those whose cumulative relationships account for the first $1/3$ of all sorted entities, and low-frequency entities are for the last $1/3$. Details are shown in Appendix B.
In practice, we first get all the hypernyms in DBpedia and classify the entities belonging to them into high-frequent and low-frequent. Then, we map the entities one-by-one to get entity pairs. The reason behind this is that not every random pair of entities can have a shared abstract question easily. For example, it’s hard to write a shared question for Einstein and Apple even though they are all high-frequency entities. But it’s easy for two specific universities to share a same abstract question. In order for our entity pairs to produce high quality abstract questions, we ensure that the two entities have the same hypernym. This allows them to have similar descriptions, making it easier to generate an abstract question.
@@ 3.2.2 Abstract Question Generation
After acquiring the entity pairs, we adopt a multi-step curation pipeline to generate high-quality abstract multiple-choice questions: (1) Generate an abstract question (without options) according to the DBpedia descriptions of both entity. (2) Separately generate the corresponding answers based on the descriptions, with length control to alleviate bias among candidate answers. (3) Generate distractors for both entities with length control. Compared to randomly selected distractors, LLM-selected distractors generation have demonstrated its effectiveness for the high relevance between distractors and designated choices Zheng et al. (2024c). (4) Formulate the final multiple-choice question using the four answer candidates above. (5) Proofread the question according to the standards below.
@@ 3.3 Expert Verification
We enlist the help of three postgraduate students, each with extensive experience in NLP research, to validate the quality of these generated questions through a sample of 200 question pairs. The instruction is the same as the standard (1) given to LLMs above. The quality of each pair is decided by majority voting. Results show that their total agreement (all 3 experts have the same judgement) is 88.0%. And 95.5% of the abstract questions are considered correct and of high quality both for the high-frequency entity and the low-frequency entity, demonstrating the reliability of our benchmark.
@@ 3.4 Main Evaluations
We experiment with a selection of LLMs on our ComparisonQA benchmark to investigate their performance on high-frequency and low-frequency questions and also the difference between them.
@@ 3.4.1 Experiment Setup
Metric: We calculate Uncertainty, Accuracy, and Macro F1-score between model predictions and ground truth labels. We apply perplexity-based uncertainty for open source LLMs and verbalized uncertainty for proprietary LLMs due to several reasons, with details explained in Appendix D. Thus, we only compare uncertainty between high and low frequency within each setting separately.
@@ 3.4.2 Results and Analysis
Evaluation results are reported in Table 1. Our observations include: (1) Huge drop in performance from high-freq to low-freq: All models suffer a performance decrease from high-frequency questions to low-frequency questions in all three settings. For instance, the accuracy and Macro F1-score of Llama-3-8B drop up to about 14 points in the few-shot setting. Proprietary models like GPT-4o are no exception. These all prove that LLM’s performance is closely related to the frequency of the knowledge in the corpus. (2) Increased Uncertainty from high-freq to low-freq: Similarly, the uncertainty of LLMs all increase from high-frequency questions to low-frequency questions. For example, the uncertainty difference of Gemma-2-9B is up to about 86 points when using zero-shot. The uncertainty in the zero-shot setting is generally higher than in the few-shot setting, and the difference is also more pronounced. We think it may be because LLMs find some familiar examples in the few-shot setting, which decreases their uncertainty. In spite of this, the difference is still clear between high-freq and low-freq. These all prove that LLMs not only perform better but also are more confident about high-frequency knowledge. (3) Few-shot helps a lot only for LLMs without instruction-tuning: Most non-instruction-tuned LLMs show a huge improvement in performance from zero-shot to few-shot, but performance is similar for those instruction-tuned LLMs. This could be because the few-shot examples only teach LLMs how to do multiple-choice questions, while those instruction-tuned ones have already learned. (4) CoT lowers uncertainty but does not awlays aid performance: It’s obvious that after the CoT inference, LLMs are more sure about their answers. However, results show that the average accuracy of GPT-4o even drops after adding CoT, which means CoT can not always help such factual questions.
@@
Llama-3 8B 75.57 61.00 68.29 14.57 65.11 (-10.47) 43.65 (-17.35) 54.38 (-13.91) ↓ 21.45 (+6.89)0
Llama-3-Instruct 8B 79.94 67.98 73.96 ↓ 11.96 68.13 (-11.82) 48.73 (-19.26) 58.43 (-15.54) ↓ 19.40 (+7.44)0
Llama-3.1 8B 74.91 62.00 68.46 ↓ 12.91 65.66 (-9.25)0 45.21 (-16.78) 55.44 (-13.02) ↓ 20.45 (+7.53)0
Llama-3.1-Instruct 8B 79.74 69.09 74.42 ↓ 10.65 72.23 (-7.51)0 55.04 (-14.06) 63.64 (-10.78) ↓ 17.20 (+6.55)0
Llama-3.2 3B 68.89 57.26 63.08 ↓ 11.63 58.89 (-10.00) 41.11 (-16.15) 50.00 (-13.07) ↓ 17.78 (+6.15)0
Llama-3.2-Instruct 3B 71.43 62.12 66.78 09.32 64.07 (-7.37)0 46.68 (-15.44) 55.37 (-11.40) ↓ 17.39 (+8.07)0
@@
Gemma-2 9B 80.10 68.36 74.23 ↓ 11.74 71.39 (-8.71)0 50.48 (-17.88) 60.94 (-13.30) 20.92 (+9.18)0
Phi-3.5-mini 4B 73.68 67.46 70.57 06.22 66.82 (-6.86)0 53.51 (-13.95) 60.16 (-10.41) ↓ 13.31 (+7.08)0
Falcon2 11B 77.11 65.92 71.52 ↓ 11.19 65.17 (-11.94) 46.88 (-19.04) 56.02 (-15.49) ↓ 18.29 (+7.10)0
Mistral-v0.3 7B 75.55 62.88 69.22 ↓ 12.67 67.86 (7.69)0 49.05 (-13.83) 58.46 (-10.76) ↓ 18.82 (+6.15)0
Mistral-v0.3-Instruct 7B 74.46 65.51 69.99 08.95 67.29 (-7.17)0 51.69 (-13.82) 59.49 (-10.50) ↓ 15.60 (+6.65)0